OpenStudy (anonymous):

Simplify the square root 50 times x to the sixth power.

OpenStudy (anonymous):

\[\sqrt{50*x ^{6}}\] That?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\sqrt{a * b}\] = \[\sqrt{a}*\sqrt{b}\] so \[\sqrt{50 * x ^{6}}\] = \[\sqrt{50}*\sqrt{x ^{6}}\]